OpenStudy (anonymous):

Can someone check if my answer is right? f(x,y) = 3x^2 - x + (10/3)y^3 - 5xy^2

OpenStudy (anonymous):

gradient? \[f(x,y) = ( 6x-1-5y ^{2}, 10y ^{2}-10xy)\]

OpenStudy (turingtest):

yes you are right!

OpenStudy (anonymous):

critical point \[6x-1-5y ^{2} = 0\] \[10y ^{2} - 10xy = 0\] so y =0 and x = 1/6

OpenStudy (turingtest):

How did you get critical points? after eliminating the y^2 you get 12x-2-10xy=0 how do you get critical points from that?

OpenStudy (anonymous):

10y^2-10xy = 0 so 10y(y-x)=0 y =0, after that i fill in the 0 in the other equation

OpenStudy (turingtest):

oh yeah... thanks

OpenStudy (anonymous):

omg.. i was asking for help xD

OpenStudy (anonymous):

but i did it correct right?

OpenStudy (turingtest):

The partial derivative part, yes. The other part makes sense, but I never had to find critical points this way before, so I can't confirm it.

OpenStudy (anonymous):

oke.. hmm i wil ask my math teacher

OpenStudy (anonymous):

thanks anyway

OpenStudy (turingtest):

Wait I remember! If you solve 10y^2-10xy = 0 for x you get x=(5y^2+1)/6 plug that into the other equation and you get -(25/3)y^3+10y^2-(5/3)y=0 (-5/3y)(5y^2-6y+1)=0 (-5/3y)(5y-1)(y-1)=0 so y={0,1/5,1} so you have to find the x's based on ALL those values

OpenStudy (turingtest):

one of those pairs is (0,1/6) find the others in the same way.

OpenStudy (anonymous):

huh i dont get it